OpenStudy (anonymous):

The equation for a circle is x2 + y2 = 25. What is the length of the circle's radius?

jimthompson5910 (jim_thompson5910):

Hint: x^2 + y^2 = r^2 is a circle with radius r

OpenStudy (whpalmer4):

Jim's way is the straightforward way, but if you like tinkering a bit, try solving the equation for y = 0 and see what points you get. Now use the distance formula or inspection to find the distance between the two points, and divide it in half. 

OpenStudy (anonymous):

is it 12.5?

OpenStudy (whpalmer4):

Mmm...no. If \(r^2 = 25\), \(r = \sqrt{25}\), right? What is the square root of 25?

OpenStudy (zzr0ck3r):

2a does not equal a^2

OpenStudy (zzr0ck3r):

ie square root of 25 is 5 not 12.5

OpenStudy (anonymous):

oh.. ok thanks (:

OpenStudy (whpalmer4):

I may have confused you with the alternate solution.

OpenStudy (anonymous):

yea kinda but still, thank you (:

OpenStudy (whpalmer4):

If we set y = 0, the equation reduces to \(x^2=25\), which is true for \(x = 5\) and \(x = -5\). The circle crosses the x-axis at (5,0) and (-5,0). Those 2 points are 10 apart, and the line between them is a diameter, not a radius, so dividing it in half, we get a radius of 5 (just as we should).